# Token-Regulated Group Relative Policy Optimization for Stable Reinforcement Learning in Large Language Models

### v1

@@ Abstract
Reinforcement learning with verifiable rewards (RLVR) has emerged as a powerful approach for strengthening the reasoning capabilities of large language models (LLMs). Among existing algorithms, Group Relative Policy Optimization (GRPO) has demonstrated strong performance, yet it suffers from a critical issue: low-probability tokens disproportionately dominate gradient updates due to their inherently large gradient magnitudes. This imbalance leads to unstable training and suppresses the contribution of high-probability tokens that are more reliable for learning. In this work, we introduce Token-Regulated Group Relative Policy Optimization (TR-GRPO), a simple yet effective extension of GRPO that assigns token-level weights positively correlated with the model’s predicted probability. By downweighting low-probability tokens and emphasizing high-probability ones, TR-GRPO mitigates gradient over-amplification while preserving informative learning signals. We provide theoretical analysis to show how token-level probability governs gradient norms which motivates our weighting design. Extensive experiments demonstrate that TR-GRPO consistently outperforms GRPO across RLVR tasks—including logic, math, and agentic reasoning—highlighting the importance of regulating token contributions during RL training and establishing TR-GRPO as a robust framework for enhancing LLM reasoning.
@@ 1 Introduction
Large language models (LLMs) have recently demonstrated remarkable progress in complex reasoning tasks such as mathematics and programming, with systems like OpenAI’s O-series (Jaech et al., 2024), DeepSeek-R1 (Guo et al., 2025), Kimi K2 (Team et al., 2025), and Qwen3 (Yang et al., 2025a) achieving state-of-the-art performance. A central driver of these advances is Reinforcement Learning with Verifiable Rewards (RLVR) (Gao et al., 2024; Lambert et al., 2024; Team et al., 2025; Guo et al., 2025; Yang et al., 2025a), which commonly employs reinforcement learning on an LLM with a rule-based outcome reward, such as a binary indicator reward for mathematical or logical validity. RLVR has emerged as a practical framework for strengthening the reasoning abilities of LLMs, offering stable and task-aligned reward signals without the need for costly human annotation or additional models. Among the RL algorithms used in this setting, Group Relative Policy Optimization (GRPO) (Shao et al., 2024) has become a popular choice due to its simplicity and strong empirical results.
Although recent progress in RLVR has been driven by new algorithms (Yu et al., 2025; Yue et al., 2025; Guan et al., 2025), cross-domain applications (Xue et al., 2025; Liu et al., 2025; Pan et al., 2025), and unexpected empirical observations  (Wang et al., 2025; Yue et al., 2025; Zhao et al., 2025), current implementations typically still treat all tokens equally without considering which ones are truly important, overlooking the fact that model responses often contain a mixture of both high-quality and low-quality tokens, especially as response length increases. Such uniform training disregards the varied functional roles tokens play in reasoning, which can hinder performance by neglecting critical tokens and amplifying noise from less relevant ones.
@@
Motivated by these observations, we ask: which tokens actually drive the update during RLVR ?. Our analysis (Theorem 3.1) shows that the gradient norm of a token scales with $1-\pi_{\theta}(o_{t})$, where $o_{t}$ is the token and $\pi_{\theta}(o_{t})$ is its model probability; hence, low-probability tokens induce larger gradients and more aggressive updates, while high-probability tokens induce smaller, more conservative updates. To check whether this aligns with semantic importance, we visualize the top 100 frequent high probability tokens and low-probability tokens ranked by their average model probability $\bar{\pi}(o_{t})=\tfrac{1}{\#\mathrm{occ}(o_{t})}\sum_{o_{t}}\pi_{\theta}(o_{t})$ (Figures 1(a)1(b)) where $\mathrm{occ}(o_{t})$ denotes the number of occurrences of token $o_{t}$. Figure 1(a) shows that high-probability tokens concentrate on mathematical and logical structure—operators, brackets, variable names, and formatting markers—where even a single error can break the entire solution, whereas low-probability tokens are mostly generic content words (e.g., “output,” “particular,” “location”) that contribute less to the core reasoning and are more easily replaced without changing meaning. This creates a clear tension: standard GRPO naturally magnifies low-probability tokens (via larger gradients) even though high-probability tokens carry the most critical signal for correctness. We therefore need a mechanism to mitigate over-updates on low-probability tokens while preserving and even reinforcing updates on high-probability ones.
@@ 2.1 Large Language Models
Most Large Language Models (LLMs) adopt a transformer decoder-only architecture (Vaswani et al., 2017), parameterized by $\theta\in\mathbb{R}^{d}$ and denoted as $\pi_{\theta}$. The basic unit of an LLM is the token, which may correspond to a word, subword, or character, and is selected from a finite vocabulary $\mathcal{V}=\{v^{1},\dots,v^{N}\}$ of size $N$. Given a prompt $q$, the model generates a sequence of tokens $o=(o_{1},\dots,o_{T})$ autoregressively. At each step $t$, the model produces a distribution over $\mathcal{V}$ conditioned on $q$ and the previously generated tokens $o_{<t}$, from which the next token is sampled:
@@ 2.2 Reinforcement Learning with Verifiable Rewards (RLVR)
Reinforcement Learning with Verifiable Rewards (RLVR)  (Gao et al., 2024; Lambert et al., 2024; Team et al., 2025; Guo et al., 2025; Yang et al., 2025a) provides a framework for optimizing LLMs in domains where correctness can be deterministically verified. Typical applications include code generation validated by unit tests, mathematical problem solving with symbolic checkers, and factual QA with exact-match rules or programmatic validators. Unlike preference-based reinforcement learning, RLVR leverages a predefined verifier to produce a reward without requiring human labels. Formally, given a dataset of prompts $\mathcal{D}$, a policy $\pi_{\theta}$, and a frozen reference model $\pi_{\mathrm{ref}}$, the objective is:
@@
where $o_{g}$ is the ground-truth answer and $\mathrm{match}(\cdot,\cdot)\in\{0,1\}$ indicates whether the generated output matches. More generally, $R_{\phi}$ may be graded (e.g., partial credit, length penalty, latency) while remaining verifiable by rule-based approaches (Wang et al., 2024) or model-based verifiers (Ma et al., 2025).
@@ 2.3 Group Relative Policy Optimization (GRPO)
Group Relative Policy Optimization (GRPO), first introduced by Shao et al. (2024), is a policy-gradient algorithm widely used for optimizing the objective in Eq. 2. Unlike the popular PPO algorithm  (Schulman et al., 2017), which requires training an additional value function alongside the LLM, GRPO removes the dependency on an explicit value model. Instead, it estimates advantages by normalizing rewards within a group of sampled responses to the same prompt. Specifically, for a prompt $q$ with $G$ sampled responses $\{o_{i}\}_{i=1}^{G}$ and associated scalar rewards $\{r_{i}\}_{i=1}^{G}$, GRPO defines a group-normalized advantage as:
@@
The effectiveness of the above normalization method can be interpreted through the lens of reward shaping: by emphasizing relative differences among candidate outputs for the same prompt, it strengthens the stability of the gradient signal and maintains its reliability, even under sparse reward conditions (Hu et al., 2020). In this study, we employ a variant of GRPO to optimize the policy model $\pi_{\theta}$. The optimization objective of GRPO can be expressed as follows:
@@ 3 Methodology
We now present our proposed approach. To see what kinds of tokens highly impact the update, we start with the derivative of the GRPO’s objective function w.r.t. the model parameters.
@@
$\Biggl[\frac{1}{\sum_{i=1}^{G}|o_{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}\underbrace{\left(\frac{\pi_{\theta}(o_{i,t})}{\pi_{\text{old}}(o_{i,t})}\hat{A}_{i,t}\cdot\mathbb{I}_{\text{trust}}\!\left(\tfrac{\pi_{\theta}(o_{i,t})}{\pi_{\text{old}}(o_{i,t})},\hat{A}_{i,t}\right)+\beta\tfrac{\pi_{\text{ref}}(o_{i,t})}{\pi_{\theta}(o_{i,t})}-\beta\right)}_{\gamma_{i,t}}\nabla_{\theta}\log\pi_{\theta}(o_{i,t})\Biggr]$ (7)
where for simplicity, we denote $\pi_{\theta}(o_{i,t}):=\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})$ and define
$\mathbb{I}_{\text{trust}}\!\left(\frac{\pi_{\theta}(o_{i,t})}{\pi_{\text{old}}(o_{i,t})},\hat{A}_{i,t}\right)=\begin{cases}0,&\text{if }\hat{A}_{i,t}>0\text{ and }\frac{\pi_{\theta}(o_{i,t})}{\pi_{\text{old}}(o_{i,t})}>1+\epsilon_{h},\\[6.0pt] 0,&\text{if }\hat{A}_{i,t}<0\text{ and }\frac{\pi_{\theta}(o_{i,t})}{\pi_{\text{old}}(o_{i,t})}<1-\epsilon_{l},\\[6.0pt] 1,&\text{otherwise}.\end{cases}$
We represent our LLM as the composition of many layers $f=f_{L}\circ f_{L-1}\circ...\circ f_{l}\circ...\circ f_{1}$ where $f$ maps from the input sequence of tokens $a_{0}$ to the output sequence of tokens $a_{L}$. Moreover, the output sequence of tokens $a_{L}$ is used to compute the logit $h_{i,t}$ for predicting the token $o_{i,t}$. Let us denote $a_{l-1}$ and $a_{l}$ as the input and output of the $l$-th layer, i.e., $a_{l}=f_{l}\left(a_{l-1},\theta_{l}\right)$ where $\theta_{l}$ represents the model parameter at the $l$-th layer. To quantify the gradient for update, we develop the following theorem.
Theorem 3.1.
Let us denote the classifier head at the output layer, the Jacobian $\frac{\partial f_{l}\left(a_{l-1},\theta_{l}\right)}{\partial a_{l-1}}$, and the gradient matrix $\frac{\partial f_{l}\left(a_{l-1},\theta_{l}\right)}{\partial\theta_{l}}$ as $W,J_{l}$, and $G_{l}$ respectively. Moreover, we further assume that $\sigma_{\textbf{min}}(W)\geq(\alpha^{W})^{2}$, $\sigma_{\textbf{max}}(W)\leq(\beta^{W})^{2}$, $\sigma_{\textbf{min}}(J_{l})\geq(\alpha^{J}_{l})^{2}$, $\sigma_{\textbf{max}}(J_{l})\leq(\beta^{J}_{l})^{2}$, and $\sigma_{\textbf{min}}(G_{l})\geq(\alpha^{G}_{l})^{2}$, $\sigma_{\textbf{max}}(G_{l})\leq(\beta^{G}_{l})^{2}$ where all lower/upper bounds are positive and $\sigma_{\text{min}}(\cdot)$, $\sigma_{\text{max}}(\cdot)$ return the smallest and the largest singular values of a matrix. We can bound the gradient norm $\|\nabla_{\theta}\log\pi_{\theta}\left(o_{i,t}\right)\|_{2}$ as
$\frac{\left(1-\pi_{\theta}\left(o_{i,t}\right)\right)}{\sqrt{L}}\sum_{l=1}^{L}\left(a^{W}a_{l}^{G}\prod_{i=l}^{L}a_{i}^{J}\right)\leq\|\nabla_{\theta}\log\pi_{\theta}\left(o_{i,t}\right)\|_{2}\leq\sqrt{2}\left(1-\pi_{\theta}\left(o_{i,t}\right)\right)\sum_{l=1}^{L}\left[b^{W}b_{l}^{G}\prod_{i=l}^{L}b_{i}^{J}\right].$ (8)
From the bounds in (8), it appears that low-probability tokens with a small $\pi_{\theta}\left(o_{i,t}\right)$ tend to have higher gradient norms $\|\nabla_{\theta}\log\pi_{\theta}\left(o_{i,t}\right)\|_{2}$, leading to more aggressive updates in Eq. (3), while high-probability tokens with a high $\pi_{\theta}\left(o_{i,t}\right)$ tend to have smaller gradient norms $\|\nabla_{\theta}\log\pi_{\theta}\left(o_{i,t}\right)\|_{2}$, leading to more humble updates.
According to the above observation, we should govern the update of low/high-probability tokens in the way that reduces the aggressiveness of low-probability token updates and maintains the consistent updates for high-probability tokens. Hinting at this motivation, we propose our approach Token-Regulated GRPO (TR-GRPO).
@@
where we have defined
$\mathbb{D}_{\mathrm{KL}}\!\left[w_{i,t};\pi_{\theta}\,\|\,\pi_{\mathrm{ref}}\right]=w_{i,t}\frac{\pi_{\mathrm{ref}}(o_{i,t}\mid q,o_{i,<t})}{\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})}-\log(w_{i,t}\frac{\pi_{\mathrm{ref}}(o_{i,t}\mid q,o_{i,<t})}{\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})})-1.$ (10)
$w_{i,t}\;=\;\operatorname{clip}\!\big(\alpha\cdot\left[\sigma\!\left(\frac{{\text{sg}[\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})}]}{\tau}\right)-\mu\right],\;L,\;U\big),$ (11)
in which $\text{sg}[\cdot]$ denotes only taking the numerical value but stopping the gradient, corresponding to the detach operation in PyTorch, $\alpha$, $\mu$ and $\tau$ are constant, while $L$ and $U$ are lower and upper clipping bounds.
To determine which tokens should be prioritized during the policy update, we compare two scenarios: (i) TR-GRPO, which assigns higher weights to high-probability tokens and lower weights to low-probability ones; and (ii) a Reverse variant that inverts TR-GRPO’s estimate by assigning 2w2-w to any token whose TR-GRPO weight is $w$, thereby giving lower weights to high-probability tokens and higher weights to low-probability ones. As shown in Figure 2, TR-GRPO (blue) exhibits steadier accuracy gains and consistently outperforms both Reverse (orange) and the GRPO baseline (green) across K&K puzzle sizes (3–7), whereas Reverse tracks GRPO closely without clear improvement. These results reinforce our motivation that emphasizing high-probability tokens leads to more stable and effective learning than over-amplifying low-probability ones.

### v2

# Sharpness-Controlled Group Relative Policy Optimization with Token-Level Probability Shaping

@@ Abstract
Reinforcement learning with verifiable rewards (RLVR) has become a practical route to improve large language model reasoning, and Group Relative Policy Optimization (GRPO) is a widely used optimizer in this setting. This paper revisits GRPO from a generalization perspective. Recent analysis shows that population performance can be controlled by a robust empirical objective that decomposes into the training loss plus a sharpness term measured by the gradient norm. We develop a token-level view of this sharpness term and show that GRPO can be dominated by a small subset of tokens with disproportionately large per-token gradients, which increases sharpness and can harm generalization. Motivated by this view, we propose Token-Regulated GRPO (TR-GRPO), which introduces a monotone probability shaping function to assign token weights based on the model’s own token probabilities, and integrates these weights into the standard GRPO. Our analysis yields a bound that isolates a probability dependent multiplicative factor in token-gradient magnitudes, explaining how probability-aware weighting suppresses sharp directions while preserving learning signal on semantically critical tokens. Experiments on logic puzzles, mathematical reasoning, and tool-augmented question answering show consistent improvements over GRPO, along with smoother gradient-norm trajectories, supporting TR-GRPO as a simple and effective generalization-oriented upgrade to GRPO for RLVR.
@@ 1 Introduction
Large language models (LLMs) have recently demonstrated remarkable progress in complex reasoning tasks such as mathematics and programming, with systems like OpenAI’s O-series (jaech2024openai), DeepSeek-R1 (guo2025deepseek), Kimi K2 (team2025kimi), and Qwen3 (yang2025qwen3) achieving state-of-the-art performance. A central driver of these advances is Reinforcement Learning with Verifiable Rewards (RLVR) (gao2024designing; lambert2024tulu; team2025kimi; guo2025deepseek; yang2025qwen3), which commonly employs reinforcement learning on an LLM with a rule-based outcome reward, such as a binary indicator reward for mathematical or logical validity. RLVR has emerged as a practical framework for strengthening the reasoning abilities of LLMs, offering stable and task-aligned reward signals without the need for costly human annotation or additional models. Among the RL algorithms used in this setting, Group Relative Policy Optimization (GRPO) (shao2024deepseekmath) has become a popular choice due to its simplicity and strong empirical results.
Although recent progress in RLVR has been driven by new algorithms (yu2025dapo; yue2025vapo; guan2025rstar), cross-domain applications (xue2025dancegrpo; liu2025flow; pan2025medvlm), and unexpected empirical observations  (wang2025reinforcement; yue2025vapo; zhao2025absolute), current implementations typically still treat all tokens equally without considering which ones are truly important, overlooking the fact that model responses often contain a mixture of both high-quality and low-quality tokens, especially as response length increases. Such uniform training disregards the varied functional roles tokens play in reasoning, which can hinder performance by neglecting critical tokens and amplifying noise from less relevant ones.
@@
Motivated by these observations, we ask: which tokens actually drive the update during RLVR ?. Under the generalization view in Theorem 3.1 and the sharpness surrogate in Eq. (8), we aim to regulate token-wise updates to reduce sharp directions while preserving learning signal on semantically critical tokens; To check whether this aligns with semantic importance, we visualize the top 100 frequent high probability tokens and low-probability tokens ranked by their average model probability $\bar{\pi}(o_{t})=\tfrac{1}{\#\mathrm{occ}(o_{t})}\sum_{o_{t}}\pi_{\theta}(o_{t})$ (Figures 1(a)1(b)) where $\mathrm{occ}(o_{t})$ denotes the number of occurrences of token $o_{t}$. Figure 1(a) shows that high-probability tokens concentrate on mathematical and logical structure—operators, brackets, variable names, and formatting markers—where even a single error can break the entire solution, whereas low-probability tokens are mostly generic content words (e.g., “output,” “particular,” “location”) that contribute less to the core reasoning and are more easily replaced without changing meaning. This creates a clear tension: standard GRPO naturally magnifies low-probability tokens (via larger gradients) even though high-probability tokens carry the most critical signal for correctness. We therefore need a mechanism to mitigate over-updates on low-probability tokens while preserving and even reinforcing updates on high-probability ones.
@@ 2.1 Large Language Models
Most Large Language Models (LLMs) adopt a transformer decoder-only architecture (vaswani2017attention), parameterized by $\theta\in\mathbb{R}^{d}$ and denoted as $\pi_{\theta}$. The basic unit of an LLM is the token, which may correspond to a word, subword, or character, and is selected from a finite vocabulary $\mathcal{V}=\{v^{1},\dots,v^{N}\}$ of size $N$. Given a prompt $q$, the model generates a sequence of tokens $o=(o_{1},\dots,o_{T})$ autoregressively. At each step $t$, the model produces a distribution over $\mathcal{V}$ conditioned on $q$ and the previously generated tokens $o_{<t}$, from which the next token is sampled:
@@ 2.2 Reinforcement Learning with Verifiable Rewards (RLVR)
Reinforcement Learning with Verifiable Rewards (RLVR)  (gao2024designing; lambert2024tulu; team2025kimi; guo2025deepseek; yang2025qwen3) provides a framework for optimizing LLMs in domains where correctness can be deterministically verified. Typical applications include code generation validated by unit tests, mathematical problem solving with symbolic checkers, and factual QA with exact-match rules or programmatic validators. Unlike preference-based reinforcement learning, RLVR leverages a predefined verifier to produce a reward without requiring human labels. Formally, given a dataset of prompts $\mathcal{D}$, a policy $\pi_{\theta}$, and a frozen reference model $\pi_{\mathrm{ref}}$, the objective is:
@@
where $o_{g}$ is the ground-truth answer and $\mathrm{match}(\cdot,\cdot)\in\{0,1\}$ indicates whether the generated output matches. More generally, $R_{\phi}$ may be graded (e.g., partial credit, length penalty, latency) while remaining verifiable by rule-based approaches (wang2024openhands) or model-based verifiers (ma2025general).
@@ 2.3 Group Relative Policy Optimization (GRPO)
Group Relative Policy Optimization (GRPO), first introduced by shao2024deepseekmath, is a policy-gradient algorithm widely used for optimizing the objective in Eq. 2. Unlike the popular PPO algorithm  (schulman2017proximal), which requires training an additional value function alongside the LLM, GRPO removes the dependency on an explicit value model. Instead, it estimates advantages by normalizing rewards within a group of sampled responses to the same prompt. Specifically, for a prompt $q$ with $G$ sampled responses $\{o_{i}\}_{i=1}^{G}$ and associated scalar rewards $\{r_{i}\}_{i=1}^{G}$, GRPO defines a group-normalized advantage as:
@@
The effectiveness of the above normalization method can be interpreted through the lens of reward shaping: by emphasizing relative differences among candidate outputs for the same prompt, it strengthens the stability of the gradient signal and maintains its reliability, even under sparse reward conditions (hu2020learning). In this study, we employ a variant of GRPO to optimize the policy model $\pi_{\theta}$. The optimization objective of GRPO can be expressed as follows:
@@ 3 Methodology
3.1 A generalization view of GRPO
GRPO-style RLVR trains a policy $\pi_{\theta}$ using samples $(q,o_{1:G})$ drawn from a joint distribution $p(q,o_{1:G})\triangleq\mathcal{Q}(q)\prod_{i=1}^{G}\pi_{\theta_{\mathrm{old}}}(o_{i}\mid q)$. In practice, optimization is performed on a finite dataset $\mathcal{S}=\{(q_{k},o_{k1:G})\}_{k=1}^{N}$, which creates a gap between the population/generalization loss $\mathcal{L}_{\mathcal{D}}(\pi_{\theta})$ (over $p$) and the empirical loss $\mathcal{L}_{\mathcal{S}}(\pi_{\theta})$ (over $\mathcal{S}$).
A key theoretical message from recent GRPO generalization analysis is that, with high probability over the sampled dataset, the population loss can be controlled by a robust version of the empirical loss:
Theorem 3.1 (Generalization via local robustness).
With probability at least $1-\delta$ over $\mathcal{S}\sim\mathcal{D}^{N}$,
$\mathcal{L}_{\mathcal{D}}(\pi_{\theta})\;\leq\;\max_{\|\theta^{\prime}-\theta\|\leq\rho}\;\mathcal{L}_{\mathcal{S}}(\pi_{\theta^{\prime}})\;+\;\text{(lower-order terms)}.$ (7)
Moreover, a first-order expansion around $\theta$ yields the approximation
$\max_{\|\theta^{\prime}-\theta\|\leq\rho}\mathcal{L}_{\mathcal{S}}(\pi_{\theta^{\prime}})\;\approx\;\mathcal{L}_{\mathcal{S}}(\pi_{\theta})\;+\;\rho\,\|\nabla_{\theta}\mathcal{L}_{\mathcal{S}}(\pi_{\theta})\|_{2},$ (8)
which highlights two complementary targets for better generalization: (i) reducing the empirical loss, and (ii) reducing the sharpness term measured by the gradient norm. This section develops a token-level view of how GRPO shapes $\|\nabla_{\theta}\mathcal{L}_{\mathcal{S}}(\pi_{\theta})\|_{2}$, and proposes a probability-aware reweighting that directly suppresses sharp directions while preserving learning signal on semantically critical tokens.
3.2 Token-level gradients in GRPO
We start from the derivative of the GRPO objective with respect to model parameters:
@@ 3.2 Token-level gradients in GRPO
$\Biggl[\frac{1}{\sum_{i=1}^{G}|o_{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}\underbrace{\left(\frac{\pi_{\theta}(o_{i,t})}{\pi_{\text{old}}(o_{i,t})}\hat{A}_{i,t}\cdot\mathbb{I}_{\text{trust}}\!\left(\tfrac{\pi_{\theta}(o_{i,t})}{\pi_{\text{old}}(o_{i,t})},\hat{A}_{i,t}\right)+\beta\tfrac{\pi_{\text{ref}}(o_{i,t})}{\pi_{\theta}(o_{i,t})}-\beta\right)}_{\gamma_{i,t}}\nabla_{\theta}\log\pi_{\theta}(o_{i,t})\Biggr].$ (9)
For simplicity, $\pi_{\theta}(o_{i,t}):=\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})$. The trust-region indicator is
$\mathbb{I}_{\text{trust}}\!\left(\frac{\pi_{\theta}(o_{i,t})}{\pi_{\text{old}}(o_{i,t})},\hat{A}_{i,t}\right)=\begin{cases}0,&\hat{A}_{i,t}>0\ \text{and}\ \frac{\pi_{\theta}(o_{i,t})}{\pi_{\text{old}}(o_{i,t})}>1+\epsilon_{h},\\[4.0pt] 0,&\hat{A}_{i,t}<0\ \text{and}\ \frac{\pi_{\theta}(o_{i,t})}{\pi_{\text{old}}(o_{i,t})}<1-\epsilon_{l},\\[4.0pt] 1,&\text{otherwise}.\end{cases}$
A crucial observation is that the sharpness term in Eq. (8) depends on the magnitude of token gradients. Under standard smoothness assumptions on transformer layers, one can bound $\|\nabla_{\theta}\log\pi_{\theta}(o_{i,t})\|_{2}$ and show it scales with $(1-\pi_{\theta}(o_{i,t}))$: low-probability tokens induce larger gradients, hence contribute more to sharpness, even when such tokens are often semantically replaceable in natural language reasoning traces.
3.3 Token-Regulated GRPO for better generalization
Motivated by Eq. (7)-(8), we aim to reduce sharpness by explicitly controlling token-level gradient magnitudes. Following the generalization development, we introduce an increasing probability shaping function $\omega(\cdot)$ and define a token weight $w_{i,t}:=\omega(\pi_{\theta}(o_{i,t}))$. This yields a weighted GRPO surrogate that remains compatible with PPO/GRPO clipping and advantage learning, while steering optimization toward flatter (less sharp) regions.
We propose Token-Regulated GRPO (TR-GRPO):
@@ 3.3 Token-Regulated GRPO for better generalization
where $r_{i,t}(\theta)\triangleq\frac{\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})}{\pi_{\theta_{\mathrm{old}}}(o_{i,t}\mid q,o_{i,<t})}$ and
$\mathbb{D}_{\mathrm{KL}}\!\left[w_{i,t};\pi_{\theta}\,\|\,\pi_{\mathrm{ref}}\right]=w_{i,t}\frac{\pi_{\mathrm{ref}}(o_{i,t}\mid q,o_{i,<t})}{\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})}-\log\!\left(w_{i,t}\frac{\pi_{\mathrm{ref}}(o_{i,t}\mid q,o_{i,<t})}{\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})}\right)-1.$ (11)
We instantiate $w_{i,t}$ with a monotone, probability-aware mapping and stop-gradient to avoid the model gaming the weights:
$w_{i,t}\;=\;\operatorname{clip}\!\Big(\alpha\cdot\big[\sigma\!\big(\tfrac{\text{sg}[\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})]}{\tau}\big)-\mu\big],\;L,\;U\Big),$ (12)
where $\text{sg}[\cdot]$ is the stop-gradient operator.
The token-weighted surrogate preserves the GRPO gradient structure up to the multiplicative factor $w_{i,t}$. Concretely, the token-wise contribution to $\nabla_{\theta}\mathcal{J}_{\text{TR-GRPO}}$ can be written as
$g_{i,t}\;=\;\gamma_{i,t}\,w_{i,t}\,\nabla_{\theta}\log\pi_{\theta}(o_{i,t}),$ (13)
with a modified trust-region indicator:
$\mathbb{I}_{\text{trust}}\!\left(\frac{\pi_{\theta}(o_{i,t})}{\pi_{\text{old}}(o_{i,t})},\hat{A}_{i,t}\right)=\begin{cases}0,&\hat{A}_{i,t}>0\ \text{and}\ \frac{\pi_{\theta}(o_{i,t})}{\pi_{\text{old}}(o_{i,t})}>w_{i,t}^{-1}(1+\epsilon_{h}),\\[4.0pt] 0,&\hat{A}_{i,t}<0\ \text{and}\ \frac{\pi_{\theta}(o_{i,t})}{\pi_{\text{old}}(o_{i,t})}<w_{i,t}^{-1}(1-\epsilon_{l}),\\[4.0pt] 1,&\text{otherwise}.\end{cases}$
Under the same layer-wise smoothness and bounded Jacobian assumptions used in standard transformer analyses, we can bound $\|g_{i,t}\|_{2}$ and isolate the key multiplicative term $w_{i,t}(1-\pi_{\theta}(o_{i,t}))$:
Theorem 3.2 (Token gradient bound for TR-GRPO).
Under the same assumptions as the standard bound on $\|\nabla_{\theta}\log\pi_{\theta}(o_{i,t})\|_{2}$, the token-gradient norm satisfies
$\frac{w_{i,t}(1-\pi_{\theta}(o_{i,t}))}{\sqrt{L}}\,|\gamma_{i,t}|\sum_{l=1}^{L}\left(a^{W}a_{l}^{G}\prod_{j=l}^{L}a_{j}^{J}\right)\leq\|g_{i,t}\|_{2}\leq\sqrt{2}\,w_{i,t}(1-\pi_{\theta}(o_{i,t}))\,|\gamma_{i,t}|\sum_{l=1}^{L}\left(b^{W}b_{l}^{G}\prod_{j=l}^{L}b_{j}^{J}\right).$ (14)
For low-probability tokens, $(1-\pi_{\theta}(o_{i,t}))$ is large while $w_{i,t}$ is small; for high-probability tokens, $(1-\pi_{\theta}(o_{i,t}))$ is small while $w_{i,t}$ is large. Thus, TR-GRPO stabilizes the product $w_{i,t}(1-\pi_{\theta}(o_{i,t}))$ across tokens, suppressing extreme token gradients that would otherwise dominate $\|\nabla_{\theta}\mathcal{L}_{\mathcal{S}}(\pi_{\theta})\|_{2}$. By Eq. (8), reducing this gradient norm directly reduces sharpness, which tightens the robust empirical bound in Eq. (7) and improves generalization.
To understand what high- vs. low-probability tokens look like in practice, we visualize the top 100 tokens by average model probability. High-probability tokens are dominated by mathematical and logical operators, brackets, variable names, and formatting markers; even small errors in these tokens can invalidate an entire solution. Low-probability tokens are mostly generic content words that contribute less to the core reasoning and are more easily replaceable. This observation motivates emphasizing high-probability tokens, since they carry the most critical signal for correctness.
To determine whether emphasizing high-probability tokens is indeed beneficial, we compare: (i) TR-GRPO, which assigns higher weights to high-probability tokens and lower weights to low-probability ones; and (ii) a Reverse variant that inverts TR-GRPO’s weight by assigning 2w2-w to any token whose TR-GRPO weight is $w$. As shown in Figure 2, TR-GRPO (blue) consistently outperforms both Reverse (orange) and the GRPO baseline (green) across K&K puzzle sizes (3–7), while Reverse tracks GRPO closely without clear improvement. Beyond accuracy, this ablation supports the generalization view: prioritizing semantically critical, high-probability tokens improves robustness of the learned behavior, whereas over-amplifying rare tokens does not.